# Uncertainty Quantification With Noise Injection in Neural Networks: A Bayesian Perspective

### v1

@@ I Introduction
Noise injection, i.e., injecting additive or multiplicative random noise to certain components in the neural network, is an effective regularization technique to enhance generalization and robustness. The targets to which noise is injected could be input [5, 6, 7], weights [8, 9, 10], and activations [11, 12]. Noise injection fundamentally introduces randomness into neural networks, naturally leading us to model uncertainty. The noise injection into the weights is especially intuitive, which is also the primary focus of this paper, as it directly introduces randomness to the parameters of the neural network. Most of the existing research on noise injection focuses on the improvement of neural network prediction performance and less on its interpretation from a Bayesian perspective. While multiplicative noise injection into activations (dropout) [2] have been studied concerning Bayesian models, noise injection into weights still lacks a Bayesian theoretical explanation.
@@ II-B Model Uncertainty Quantification
BNNs [1] replace deterministic weights in traditional neural networks with distribution, allowing for the estimation of the model uncertainty. Due to the intractability of the posterior distribution, various MCMC-based and variational inference-based methods are proposed to train BNNs, including Bayes by Backprop [14] (BBB), Stochastic Gradient Langevin Dynamics [15] (SGLD), Stochastic Gradient MCMC [16] (SG-MCMC), etc.
@@ III-B Weight Noise Injection As a Bayesian Approximation
We theoretically prove that weight noise injection serves as a Bayesian approximation to the deep Gaussian process model. Let $\hat{\textbf{y}}$ be the output and x be the input of a neural network respectively, with dimensions $D$ and $Q$. We start the discussion from a single-layer NN. Assuming there are $K$ units in the hidden layer. Let $\textbf{W}_{1}$ be the weight matrix of dimensions $Q\times K$, connecting the input layer and the hidden layer, and $\textbf{W}_{2}$ be the weight matrix of dimensions $K\times D$, connecting the hidden layer and the output layer. Let b be a $K$ dimensional vector, and $\sigma(\cdot)$ is the element-wise non-linearity such as ReLU or tanh. Thus, a standard neural network can be represented as:
@@
where $l(\hat{\textbf{y}},\textbf{y})$ is the mean square error for regression tasks and cross entropy for classification tasks. Herein, we only discuss regression tasks, however, the extension to classification tasks can be easily achieved.
We are able to show that minimizing $\mathcal{L}_{\text{NN-NI}}$ in (3) is equivalent to maximizing the ELBO of a particular Gaussian process. For the Gaussian process in (III-A1), define the covariance function $\textbf{K}(\textbf{x},\textbf{y})=\int p(\textbf{w})p(b)\sigma(\textbf{w}^{T}\textbf{x}+b)\sigma(\textbf{w}^{T}\textbf{y}+b)\text{d}\textbf{w}\text{d}\textbf{b}$ with $p(\textbf{w})$ a standard multivariate normal distribution of dimensionality $Q$ and some distribution $p(b)$. Using Monte Carlo integration with $K$ terms to approximate the integral above: $\hat{\textbf{K}}(\textbf{x},\textbf{y})=\frac{1}{K}\sum_{k=1}^{K}\sigma(\textbf{w}_{k}^{T}\textbf{x}+b_{k})\sigma(\textbf{w}_{k}^{T}\textbf{y}+b_{k})$ with $_{k}\sim p(\textbf{w})$ w ) and $b_{k}\sim p(b)$, where $K$ is the number of hidden units in our single hidden layer NN approximation, and replacing the variational distribution $q(\textbf{W}_{1},\textbf{W}_{2},\textbf{b}):=q(\textbf{W}_{1})q(\textbf{W}_{2})q(\textbf{b})$:
@@
$q(\boldsymbol{b})$ $=\mathcal{N}(\boldsymbol{m},\boldsymbol{\sigma}^{2}\textbf{I}_{K}),$
@@
where $\tau$ is the model precision and $\hat{\textbf{y}}_{n}$ is the output of a NN with weight noise injection, which is defined in (2). Details of the proof are provided in supplemental materials. Therefore, minimizing $\mathcal{L}_{\text{GP-MC}}$ is equivalent to minimizing $\mathcal{L}_{\text{NN-NI}}$ in (3). The above proof can be generalized to multi-layer neural networks. Thus, a neural network with weight noise injection could be approximated by a GP model, and according to Theorem 1, GP models have equivalence to BNNs. Accordingly, the predictive posterior distribution is:
@@
In practical terms, this entails conducting T forward passes during inference, computing the mean of the predicted values as the prediction, and calculating the variance of the predicted values as an estimate of uncertainty. We name it Monte Carlo Noise Injection (MCNI). In the following, we will refer to the methods where the noise level $\alpha$ is set as a fixed value or treated as a learnable parameter as MCNI (fixed) and MCNI (learned), respectively.
@@ IV-B Metrics
We use Prediction Interval Coverage Probability (PICP) and Mean Prediction Interval Width (MPIW) as metrics to evaluate the performance of the model and compare it with MC Dropout as the baseline. The formulas of PCIP and MPIW are as follows:
@@
where $\hat{y}_{i}^{\text{U}}$ and $\hat{y}_{i}^{\text{L}}$ are the upper and lower bounds of the prediction interval for the $i$-th data points, and $N$ is the number of samples. Let $\hat{y}$ be the predicted mean and $\hat{\sigma}_{y}$ be the predicted standard deviation. We define the upper and lower bounds of the prediction interval as $\hat{y}^{U}=\hat{y}+3\hat{\sigma}_{y}$ and $\hat{y}^{L}=\hat{y}-3\hat{\sigma}_{y}$, respectively.
@@ V-A Datasets and Models
We use 8 classical regression datasets from the UCI Machine Learning Repository. We utilize a two-layer neural network with 50 units, except for the protein structure dataset, where we employ a two-layer neural network with 100 units. With a batch size of 32, an Adam optimizer with a momentum of 0.9 is used. The maximum number of epochs is 2000 and an early stopping with patience = 3 is used. We conduct a grid search for the other hyperparameters, including learning rate, weight decay, dropout probability for MC dropout, noise level for MCNI (fixed), and initial value of learnable noise level for MCNI (learned) as listed in Table I. We employ cross-validation to identify the set of parameters that yields the minimum validation loss for each model.
@@ V-B Metrics
For regression tasks, we use root mean squared error (RMSE) and mean standardized log loss (MSLL). MSLL is the difference from the negative log-likelihood ($-\log{p(\hat{y}|X,y,\hat{x})}=\frac{1}{2}\log(2\pi\hat{\sigma}_{y}^{2})+\frac{(y-\hat{y})^{2}}{2\hat{\sigma}_{y}^{2}}$) of the baseline model. For classification tasks, we use the accuracy, the expected calibration error (ECE) [28], and the Brier scor [29]. Among these, RMSE and accuracy measure predictive performance, while MSLL, ECE, and Brier score assess calibration performance, which are crucial indicators of uncertainty quality.
@@ V-C Demonstration of Uncertainty on CIFAR10
We select an image labeled as class 8 (ship) from CIFAR10 as an example, perform 100 forward passes, and plot the softmax outputs as shown in Fig. LABEL:softmax. To compare the robustness of different models, we gradually add Gaussian noise with increasing variance to the image. As the noise increases, the model misclassifies the image as class 9 (truck). It can be observed that for the vanilla model, when the standard deviation of the noise reaches or exceeds 0.06, the model begins to misclassify. For MC Dropout and MCNI, as the noise increases, the softmax outputs become more dispersed, indicating higher uncertainty. For MC Dropout and MCNI (fixed), when the standard deviation of the input noise is greater than or equal to 0.07, the softmax output for class 9 exceeds that of class 8. However, the uncertainty in the softmax output of MCNI (fixed) is higher than MC dropout. This suggests that our model is better at signaling a lack of confidence in its predictions. MCNI (learned) appears more robust to noise, as it makes correct predictions until the noise standard deviation reaches 0.1.
@@ V-D Results
TABLE II: Test RMSE and MSLL of the deterministic NN, MC dropout, and MCNI on the UCI regression datasets.
Dataset Test RMSE Test MSLL
MC dropout MCNI (fixed) MCNI (learned) Deterministic NN MC dropout MCNI (fixed) MCNI (learned)
Boston Housing 0.1308 0.1308 0.1377 0.1298 0 -15.5324 -66.1618
Concrete Strength 0.0996 0.0996 0.1001 0.0987 0 -2.1656 -43.237
Energy Efficiency 0.0102 0.0102 0.0078 0.0097 0 12.8721 22.2481
Kin8nm 0.0757 0.0757 0.0716 0.0738 0 -43.7374 -22.4681
Power Plant 0.0498 0.0498 0.0491 0.0496 0 -461.9055 -221.2814
Protein Structure 0.4762 0.4762 0.4458 0.4637 0 -558.3141 -952.2809
Wine Quality Red 0.5974 0.5974 0.5913 0.5961 0 14.6659 11.8569
Yacht 0.0406 0.0406 0.0398 0.0377 0 -20.0878 -20.2547
TABLE III: Test accuracy, ECE and Brier score of the deterministic NN, MC dropout, and MCNI on CIFAR10. The experiment is repeated five times and the standard deviation errors are calculated.
Model Accuracy ECE Brier score
Deterministic $\mathbf{72.05\pm 0.287}$ $0.0198\pm 0.003$ $0.3874\pm 0.003$
MC dropout $70.79\pm 0.706$ $0.0187\pm 0.004$ $0.3885\pm 0.010$
MCNI(fixed) $71.08\pm 0.793$ $0.0192\pm 0.005$ $0.3854\pm 0.010$
MCNI(learned) $71.19\pm 0.802$ $\mathbf{0.0143\pm 0.005}$ $\mathbf{0.3833\pm 0.011}$
@@ VI-A Inference Time Analysis
Inference is performed on the CIFAR10 test set using the same hardware and model as described in V-A, with a batch size of 500. The inference time is recorded, and the experiment is repeated 10 times to obtain the average result, which is presented in Table IV. It can be observed that there is no significant difference in the time taken for MCNI and MC dropout. Furthermore, the increase in time for 10 forward propagations, compared to a single forward propagation, is minimal.
@@ VI-B Conclusion
In this paper, we establish theoretically that introducing noise into a neural network’s weights is equivalent to Bayesian inference on a deep Gaussian process, by selecting a specific covariance function and variational distributions. Consequently, we propose a Monte Carlo Noise Injection (MCNI) technique, which involves incorporating noise into the parameters during training and executing forward propagation multiple times during inference to evaluate the uncertainty of prediction. Our assessment of the predictive distribution using a toy dataset reveals that our method produces better results in comparison to MC dropout. Moreover, we evaluate predictive performance, calibration performance and selective performance on both regression and classification tasks using UCI regression datasets and CIFAR10, demonstrating that our method outperforms the baseline model.
Since weight noise injection is commonly used in adversarial training and model uncertainty can also help identify adversarial samples, a potential future research direction could involve integrating the two. Additionally, exploring the application of this method in other domains, such as uncertainty quantification in reinforcement learning, could be considered. Our future research will address also non-Gaussian noise injection motivated by the results in [7] and [30]. Further, we will also study noise injection in graphs and consider adaptive signal processing methods for noise robustness [31, 32]. We will also study the effect of NN architecture on the noise injection performance [33].

### v2

@@ I Introduction
Noise injection, i.e., injecting additive or multiplicative random noise to certain components in the neural network, is an effective regularization technique to enhance generalization and robustness. The targets to which noise is injected could be input [5, 6, 7], weights [8, 9, 10], and activations [11, 12]. Noise injection fundamentally introduces randomness into neural networks, naturally leading us to model uncertainty. The noise injection into the weights is especially intuitive, which is also the primary focus of this paper, as it directly introduces randomness to the parameters of the neural network. Most of the existing research on noise injection focuses on the improvement of neural network prediction performance and less on its interpretation from a Bayesian perspective. While multiplicative noise injection into activations (dropout) [2] has been studied concerning Bayesian models, noise injection into weights still lacks a Bayesian theoretical explanation.
@@ II-B Model Uncertainty Quantification
BNNs [1] replace deterministic weights in traditional neural networks with distributions, allowing for the estimation of the model uncertainty. Due to the intractability of the posterior distribution, various MCMC-based and variational inference-based methods have been proposed to train BNNs, including Bayes by Backprop [14] (BBB), Stochastic Gradient Langevin Dynamics [15] (SGLD), Stochastic Gradient MCMC [16] (SG-MCMC), etc.
@@ III-B Weight Noise Injection As a Bayesian Approximation
We theoretically prove that weight noise injection serves as a Bayesian approximation to the deep Gaussian process model. Let $\hat{\textbf{y}}$ be the output and x be the input of a neural network respectively, with dimensions $D$ and $Q$. We start the discussion from a single-layer NN. Assuming there are $K$ units in the hidden layer. Let $\textbf{W}_{1}$ be the weight matrix of dimensions $Q\times K$, connecting the input layer and the hidden layer, and $\textbf{W}_{2}$ be the weight matrix of dimensions $K\times D$, connecting the hidden layer and the output layer. Let b be a $K$ dimensional vector, and $\sigma(\cdot)$ be the element-wise non-linearity such as ReLU or tanh. Thus, a standard neural network can be represented as:
@@
where $l(\hat{\textbf{y}},\textbf{y})$ is the mean squared error for regression tasks and cross entropy for classification tasks. Herein, we only discuss regression tasks, however, the extension to classification tasks can be easily achieved.
We are able to show that minimizing $\mathcal{L}_{\text{NN-NI}}$ in (3) is equivalent to maximizing the ELBO of a particular Gaussian process. For the Gaussian process in (III-A1), define the covariance function $\textbf{K}(\textbf{x},\textbf{y})=\int p(\textbf{w})p(b)\sigma(\textbf{w}^{T}\textbf{x}+b)\sigma(\textbf{w}^{T}\textbf{y}+b)\text{d}\textbf{w}\text{d}\textbf{b}$ with $p(\textbf{w})$ as a standard multivariate normal distribution of dimensionality $Q$ and some distribution $p(b)$. Using Monte Carlo integration with $K$ terms to approximate the integral above: $\hat{\textbf{K}}(\textbf{x},\textbf{y})=\frac{1}{K}\sum_{k=1}^{K}\sigma(\textbf{w}_{k}^{T}\textbf{x}+b_{k})\sigma(\textbf{w}_{k}^{T}\textbf{y}+b_{k})$ with $_{k}\sim p(\textbf{w})$ w ) and $b_{k}\sim p(b)$, where $K$ is the number of hidden units in our single hidden layer NN approximation, and replacing the variational distribution $q(\textbf{W}_{1},\textbf{W}_{2},\textbf{b}):=q(\textbf{W}_{1})q(\textbf{W}_{2})q(\textbf{b})$:
@@
$q(\boldsymbol{b})$ $=\mathcal{N}(\textbf{m},\boldsymbol{\sigma}^{2}\textbf{I}_{K}),$
@@
where $\textbf{M}_{1}=[\textbf{m}_{q}]_{q=1}^{Q}$, $\textbf{M}_{2}=[\textbf{m}_{k}]_{k=1}^{K}$, $N$ is the number of inputs, $\tau$ is the model precision and $\hat{\textbf{y}}_{n}$ is the output of a NN with weight noise injection, which is defined in (2). Details of the proof are provided in supplemental materials. Therefore, minimizing $\mathcal{L}_{\text{GP-MC}}$ is equivalent to minimizing $\mathcal{L}_{\text{NN-NI}}$ in (3). The above proof can be generalized to multi-layer neural networks, which is shown in supplemental materials. Thus, a neural network with weight noise injection could be approximated by a GP model, and according to Theorem 1, GP models have equivalence to BNNs. Accordingly, the predictive posterior distribution is:
@@
In practical terms, this entails conducting T forward passes during inference, computing the mean of the predicted values as the prediction, and calculating the variance of the predicted values as an estimate of uncertainty. We name it Monte Carlo Noise Injection (MCNI). In the following, we will refer to the methods where the noise level $\alpha$ is set to a fixed value or treated as a learnable parameter as MCNI (fixed) and MCNI (learned), respectively.
@@ IV-B Metrics
We use Prediction Interval Coverage Probability (PICP) and Mean Prediction Interval Width (MPIW) as metrics to evaluate the performance of the model and compare it with MC Dropout as the baseline. The formulas for PICP and MPIW are as follows:
@@
where $\hat{y}_{i}^{\text{U}}$ and $\hat{y}_{i}^{\text{L}}$ are the upper and lower bounds of the prediction interval for the $i$-th data point, and $N$ is the number of samples. Let $\hat{y}$ be the predicted mean and $\hat{\sigma}_{y}$ be the predicted standard deviation. We define the upper and lower bounds of the prediction interval as $\hat{y}^{U}=\hat{y}+3\hat{\sigma}_{y}$ and $\hat{y}^{L}=\hat{y}-3\hat{\sigma}_{y}$, respectively.
@@ V-A Datasets and Models
We use 8 classical regression datasets from the UCI Machine Learning Repository. For the choice of hyperparameters we refer to [27]. Specifically, we utilize a two-layer neural network with 50 units, except for the protein structure dataset, where we employ a two-layer neural network with 100 units. With a batch size of 32, an Adam optimizer with a momentum of 0.9 is used. The maximum number of epochs is 2000 and an early stopping with patience = 3 is used. We conduct a grid search for the other hyperparameters, including learning rate, weight decay, dropout probability for MC dropout, noise level for MCNI (fixed), and initial value of learnable noise level for MCNI (learned) as listed in Table I. We employ cross-validation to identify the set of parameters that yields the minimum validation loss for each model.
@@ V-B Metrics
For regression tasks, we use root mean squared error (RMSE) and mean standardized log loss (MSLL). MSLL is defined as the difference from the negative log-likelihood ($-\log{p(\hat{y}|X,y,\hat{x})}=\frac{1}{2}\log(2\pi\hat{\sigma}_{y}^{2})+\frac{(y-\hat{y})^{2}}{2\hat{\sigma}_{y}^{2}}$) of the baseline model (MC dropout). Therefore, the MSLL of MC dropout is always 0. For classification tasks, we use the accuracy, the expected calibration error (ECE) [29], and the Brier score [30]. Among these, RMSE and accuracy measure predictive performance, while MSLL, ECE, and Brier score assess calibration performance, which are crucial indicators of uncertainty quality.
@@
TABLE II: Test RMSE and MSLL (mean $\pm$ std) of the deterministic NN, MC dropout, and MCNI on the UCI regression datasets. The experiment is repeated five times.
Dataset Test RMSE Test MSLL
Deterministic NN MC dropout MCNI (fixed) MCNI (learned) MC dropout MCNI (fixed) MCNI (learned)
Boston Housing $0.1308\pm 0.0058$ $0.1340\pm 0.0068$ $0.1377\pm 0.0105$ $\boldsymbol{0.1298}\pm 0.0067$ $0\pm 0.86$ $-15.53\pm 0.74$ $\boldsymbol{-66.16}\pm 0.90$
Concrete Strength $0.0996\pm 0.0063$ $0.1003\pm 0.0075$ $0.1001\pm 0.0068$ $\boldsymbol{0.0987}\pm 0.0066$ $0\pm 0.53$ $-2.17\pm 0.66$ $\boldsymbol{-43.24}\pm 0.64$
Energy Efficiency $0.0102\pm 0.0019$ $0.0090\pm 0.0030$ $\boldsymbol{0.0078}\pm 0.0023$ $0.0097\pm 0.0026$ $\boldsymbol{0}\pm 0.19$ $12.87\pm 0.21$ $22.25\pm 0.16$
Kin8nm $0.0757\pm 0.0011$ $0.0752\pm 0.0018$ $\boldsymbol{0.0716}\pm 0.0013$ $0.0738\pm 0.0013$ $0\pm 1.34$ $\boldsymbol{-43.74}\pm 1.22$ $-22.47\pm 1.43$
Power Plant $0.0498\pm 0.0003$ $0.0496\pm 0.0006$ $\boldsymbol{0.0491}\pm 0.0006$ $0.0496\pm 0.0008$ $0\pm 2.32$ $\boldsymbol{-461.91}\pm 2.43$ $-221.28\pm 2.33$
Protein Structure $0.4762\pm 0.0123$ $0.4649\pm 0.0146$ $\boldsymbol{0.4458}\pm 0.0139$ $0.4637\pm 0.0140$ $0\pm 1.36$ $-558.31\pm 1.02$ $\boldsymbol{-952.28}\pm 1.86$
Wine Quality Red $0.5974\pm 0.0108$ $0.5968\pm 0.0076$ $\boldsymbol{0.5913}\pm 0.0075$ $0.5961\pm 0.0068$ $\boldsymbol{0}\pm 1.66$ $14.67\pm 1.46$ $11.86\pm 1.53$
Yacht $0.0406\pm 0.0010$ $0.0388\pm 0.0017$ $0.0398\pm 0.0018$ $\boldsymbol{0.0377}\pm 0.0018$ $0\pm 1.20$ $-20.09\pm 1.32$ $\boldsymbol{-20.25}\pm 1.24$
@@ V-C Demonstration of Uncertainty on CIFAR10
We select an image labeled as class 8 (ship) from CIFAR10 as an example, perform 100 forward passes, and plot the softmax outputs as shown in Fig. LABEL:softmax. To compare the robustness of different models, we gradually add Gaussian noise with increasing variance to the image. As the noise increases, the model misclassifies the image as class 9 (truck). It can be observed that for the vanilla model, when the standard deviation of the noise reaches or exceeds 0.06, the model begins to misclassify. For MC Dropout and MCNI, as the noise increases, the softmax outputs become more dispersed, indicating higher uncertainty. For MC Dropout and MCNI (fixed), when the standard deviation of the input noise is greater than or equal to 0.07, the softmax output for class 9 exceeds that of class 8. However, the uncertainty in the softmax output of MCNI (fixed) is higher than that of MC dropout. This suggests that our model is better at signaling a lack of confidence in its predictions. MCNI (learned) appears more robust to noise, as it makes correct predictions until the noise standard deviation reaches 0.1.
TABLE III: Test accuracy, ECE and Brier score (mean $\pm$ std) of the deterministic NN, MC dropout, and MCNI on CIFAR10. The experiment is repeated five times.
Model Accuracy ECE Brier score
Deterministic $\mathbf{72.05\pm 0.29}$ $0.0198\pm 0.003$ $0.3874\pm 0.003$
MC dropout $70.79\pm 0.71$ $0.0187\pm 0.004$ $0.3885\pm 0.010$
MCNI(fixed) $71.08\pm 0.79$ $0.0192\pm 0.005$ $0.3854\pm 0.010$
MCNI(learned) $71.19\pm 0.80$ $\mathbf{0.0143\pm 0.005}$ $\mathbf{0.3833\pm 0.011}$
@@ VI-A Inference Time Analysis
Inference is performed on the CIFAR10 test set using the same settings as described in V-A, with a batch size of 500. The inference time is recorded, and the experiment is repeated 10 times to obtain the average result, which is presented in Table IV. It can be observed that there is no significant difference in the time taken for MCNI and MC dropout. Furthermore, the increase in time for 10 forward propagations, compared to a single forward propagation, is minimal.
@@ VI-B Conclusion
In this paper, we establish theoretically that introducing noise into a neural network’s weights is equivalent to Bayesian inference on a deep Gaussian process, by selecting a specific covariance function and variational distributions. Consequently, we propose a Monte Carlo Noise Injection (MCNI) technique, which involves incorporating noise into the parameters during training and executing forward propagation multiple times during inference to evaluate the uncertainty of prediction. Our assessment of the predictive distribution using a toy dataset reveals that our method produces better results compared to MC dropout. Moreover, we evaluate predictive performance, calibration performance and selective performance on both regression and classification tasks using UCI regression datasets and CIFAR10, demonstrating that our method outperforms the baseline model.
Since weight noise injection is commonly used in adversarial training and model uncertainty can also help identify adversarial samples, a potential future research direction could involve integrating the two. Exploring the application of this method in other domains, such as uncertainty quantification in reinforcement learning, could be considered. Our future research will also address non-Gaussian noise injection motivated by the results in [7]. Further, we will also study noise injection in graphs and consider adaptive signal processing methods for noise robustness [31, 32], and the effect of NN architecture on the noise injection performance [33].